You purchased a machine for $1.00 million three years ago and have been applying​ straight-line depreciation to zero for a​ seven-year life. Your tax rate is 35%. If you sell the machine today​ (after three years of​ depreciation) for $700,000​, what is your incremental cash flow from selling the​ machine? Your total incremental cash flow will be. (Round to the nearest​ cent.)

Please show all steps needed to reach final answer
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(1,000,000/7)=$142857.143

Hence book value as on date of sale=Cost-Accumulated depreciation

=1,000,000-(142857.143*3)

=571428.571

Hence gain on sale=(700,000-571428.571)=128571.429

Hence incremental cash flow=Sale proceeds-(Tax rate*Gain on sale)

=700,000-(128571.429*0.35)

=$655,000